Three pairs of shoes are displayed in a row. In how many ways can the shoes be arranged
(a) with no restrictions on their positions,
(b) if mates must be kept together but mates can be placed in either order,
(c) if mates must be kept together and left and right mates are kept in order, left shoe then right?

Answer by TimothyLamb(4379)

Three pairs of shoes are displayed in a row. In how many ways can the shoes be arranged:
---
(a) with no restrictions on their positions:
6*5*4*3*2*1 = 720
---
(b) if mates must be kept together but mates can be placed in either order:
6*1*4*1*2*1 = 48
---
(c) if mates must be kept together and left and right mates are kept in order, left shoe then right:
3*1*2*1*1*1 = 6
---
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php